# Non-Intrusive Graph-Based Bot Detection for E-Commerce Using Inductive Graph Neural Networks

### v4

@@ III-B Feature Design
Content (URL) node features: Each URL node has attributes describing page context, including:
  • Type/context: page category (e.g., product, search, home).
  • Global statistics: popularity/rarity and coarse sensitivity tags for special endpoints.
These URL attributes contextualize access patterns (e.g., concentrated browsing of rare or sensitive endpoints).
@@ IV-A Dataset and Experimental Setup
Dataset: We use anonymized server logs from a representative mid-sized e-commerce platform over two weeks, built via a hybrid (semi-synthetic) strategy for high-confidence labels. Background traffic consists of real production sessions ($\sim$80K) after removing sessions with <2<2 requests and truncating extreme outliers. Bots constitute a small fraction ($\sim$5%) and are drawn from verified real-world attacks (honeypots/trap URLs) as well as controlled injections (scrapers/headless browsers). The session–URL graph has on the order of $10^{5}$ edges (tens of thousands of sessions; thousands of URLs) with a power-law degree distribution; identifiers are anonymized/hashed. We use 10% validation and 10% test splits with similar class proportions, and a chronologically later test split to emulate deployment and evaluate inductive generalization.
@@ IV-B Performance Comparison
Note: indicates results that align with the qualitative metrics reported in the deployment analysis.

### v5

@@ III-B Feature Design
Content (URL) node features: Content (URL) node features are intentionally designed to be coarse-grained and privacy-preserving. We do not use raw URL strings, path tokens, query parameters, or any user-generated content (e.g., search terms or identifiers) as model inputs. Instead, each URL is mapped to a small set of high-level semantic attributes, including page category (e.g., product, category, search, checkout) and global access statistics such as relative popularity or rarity.
All URL identifiers are anonymized via one-way hashing prior to graph construction, and the model operates exclusively on these abstracted features. This design follows data minimization principles and avoids exposure to personally identifiable information (PII), enabling non-intrusive deployment without client-side instrumentation or explicit user consent.
  • Type/context: page category (e.g., product, category, search, checkout).
  • Global statistics: relative popularity/rarity and coarse sensitivity tags for special endpoints.
These URL attributes contextualize access patterns (e.g., concentrated browsing of rare or sensitive endpoints) while preserving privacy.
@@ IV-A Dataset and Experimental Setup
Dataset: We use anonymized server logs from a representative mid-sized e-commerce platform over two weeks, built via a hybrid (semi-synthetic) strategy for high-confidence labels. Background traffic consists of real production sessions ($\sim$80K) after removing sessions with <2<2 requests and truncating extreme outliers. Bots constitute a small fraction ($\sim$5%) and are drawn from verified real-world attacks (honeypots/trap URLs) as well as controlled injections (scrapers/headless browsers). The session–URL graph has on the order of $10^{5}$ edges (tens of thousands of sessions; thousands of URLs) with a power-law degree distribution. All session and URL identifiers are anonymized using one-way hashing, and no raw URLs, query parameters, or user-specific identifiers are retained in the dataset. We use 10% validation and 10% test splits with similar class proportions, and a chronologically later test split to emulate deployment and evaluate inductive generalization.
@@ IV-D Cold-Start Simulation Experiment
IV-E Robustness to Distribution Shift and Unseen Targets
To further characterize the inductive challenge beyond the Week 1 $\rightarrow$ Week 2 cold-start split, we quantify distribution shift between weeks and evaluate an extreme subset that emphasizes previously unseen targets.
1) Quantifying distribution shift. We measure shift in both behavioral statistics and content usage. Specifically, we compute Jensen–Shannon (JS) divergence between Week 1 and Week 2 session-level distributions (request rate and session duration), obtaining a value of 0.083, indicating a non-trivial drift in behavioral patterns. Structurally, 19.2% of the URL nodes appearing in the Week 2 graph were unseen during Week 1, confirming that the test period is not a near-i.i.d. continuation of training and strictly requires inductive generalization.
2) Extreme case: unseen-target sessions. To investigate more extreme inductive conditions, we construct a hard subset from Week 2 consisting of sessions whose visited URLs are entirely unseen in Week 1 (i.e., all session–URL edges connect to URL nodes absent from the training graph). This setting simulates attacks that target newly launched inventory or novel page categories that were not present during training. In our data, this subset contains 1,428 sessions.
We evaluate both GraphSAGE and the session-only MLP on this subset using the same operating procedure as the main experiments. As summarized in Table 3, performance degrades for both models, as expected under reduced neighborhood overlap and limited historical context. However, GraphSAGE exhibits a substantially smaller drop than the MLP baseline and maintains a clear advantage. This indicates that GraphSAGE’s gains stem from learning generalizable interaction patterns via feature-driven aggregation (including coarse URL semantic attributes such as page category and sensitivity tags), rather than memorizing specific “bad” nodes.
3) Static graph modeling vs. temporal methods. While session behaviors are inherently time-ordered, our objective is not trajectory prediction but malicious session identification under sparse and evolving data. In this setting, explicit temporal modeling is not always advantageous. Many bot sessions are short-lived, incomplete, or intentionally obfuscated, where fine-grained temporal dependencies are weak or noisy. By encoding temporal information implicitly through session structure (e.g., URL transitions) and node attributes, a static graph formulation provides a robust and computationally efficient representation. More complex temporal graph models such as TGN or trajectory-based Transformers may offer benefits in settings with dense, long-horizon user trajectories, which we leave as future work.
Table 3: Performance under extreme cold-start on unseen-target sessions (Week 2 subset).
Model Week 2 Overall AUC Unseen-Target Subset AUC Drop
Session-level MLP (Baseline) 0.8500 0.7210 -15.2%
GraphSAGE (Ours) 0.9630 0.8890 -7.7%
IV-F Impact of Session Length and Graph Sparsity
To further understand the operational boundaries of graph-based detection under sparse interaction regimes, we analyze model performance across different session lengths.
In our data, the majority of sessions fall within the short to medium range (3–50 URL visits), while extremely short (0–2) and very long (>>50) sessions are less frequent. For very short sessions, performance degrades for all models due to insufficient relational context, and the advantage of graph-based aggregation is limited. Performance improves substantially for short and medium-length sessions, where sufficient interaction history enables effective message passing and feature aggregation. For very long sessions, performance slightly decreases, likely due to increased noise and repetitive navigation patterns, yet GraphSAGE consistently maintains an advantage over the session-only baseline.
Overall, this analysis indicates a practical minimum interaction threshold of approximately three URL visits for reliable classification and clarifies the session-length regimes in which the proposed method is most effective.
@@ VI-A Future Work
Appendix A Supplementary Material
The overall values reported here summarize performance within the stratified evaluation context and are not directly comparable to the aggregate AUC reported in Table 1.
Table 4: Stratified performance by session length (supplementary).
[…]
Very short (0–2) 0.6200 0.6640 +0.0440
Short (3–10) 0.7800 0.8880 +0.1080
Medium (11–50) 0.8200 0.9430 +0.1230
Long (>50>50) 0.7600 0.9000 +0.1400
Overall 0.8100 0.9300 +0.1200